OpenStudy (anonymous):

Select the best possible first step to solving the following system by first eliminating the x variable. 5x – 7y = –4 3x – 10y = –14

OpenStudy (anonymous):

First you should multiply both equations so that both get 15x (the first equation by 3 and the second by 5) in order to be able to eliminate the x-variable, you should get: \[15x-21y=-12\]\[15x-50y=-70\] Then define 15x: \[15x=21y-12 \] or \[15x=50y-70\] Then simply substitute one of the 15 x with their 21y-12 or 50y-70. \[21y-12=50y-70\] \[-29y=-58\] This should give you: x=y=2